Let f be a non- negative measurable function. Then there is a sequence<pi(n)> of simple function with pi(n+1)à ƒ ƒ ¢ ‰ ¥pi(n),such that f=lim pi(n) at each point of x.If f is defined on a Sigma- finite measure space.we may choose the function pi(n) so that each vanishes outside a set of a finite measure
1
Expert's answer
2021-10-20T13:07:05-0400

For each k ≥ 1, we partition the interval [0,∞] into disjoint interval as follows:

[0,∞] = [0, 1) "\bigcup" [1, 2)"\bigcup" · · · [n − 1, n) "\bigcup" [n,∞]


"\bigcup^{n2^n}_{ i=1}" "[\frac{i-1}{2^n},\frac{i}{2^n})\bigcup [n,\infin)"


Define, for each 1 ≤ i ≤ n2n ,

Eni = f −1"([\frac{i-1}{2^n},\frac{i}{2^n}))"


and En = f −1 ([n, ∞]). Define the function "\pi_n" (x) as follows:


"\displaystyle{\sum_{i=1}^{n2^n}}\frac{i-1}{2^n}\chi _{E_{n,i}}+n\chi_{E_n}"


If f(x) < ∞, then there exists a positive integer n such that f(x) < n. Since x ∈ En,i for some

1 ≤ i ≤n2n , we have

0 ≤ f(x) −"\pi_n"(x) ≤ "\frac{i}{2^n}-\frac{i-1}{2^n}=\frac{1}{2^n}"


If f(x) = ∞, then "\pi_n"(x) = n for every n. Thus, for each x ∈ E,

"\displaystyle{\lim_{n\to\infin}}\pi_n(x)=f(x)"


This gives us a sequence "\pi_n"(x) of simple functions on E converging to f(x).

We now show that "\pi_n"(x) ≤ "\pi_{n+1}"(x) for every integer x ∈ E. To do so, notice that for each

1 ≤ i ≤ "n2^n" ,


"E_{n+1,2i-1}\bigcup E_{n+1,2i}=f^{-1}([\frac{2i-2}{2^{n+1}},\frac{2i-1}{2^{n+1}}))\bigcup f^{-1}([\frac{2i-1}{2^{n+1}},\frac{2i}{2^{n+1}}))=f^{-1}([\frac{i-1}{2^{n}},\frac{i}{2^{n}}))=E_{ni}"


If x ∈ En,i for some 1 ≤ i ≤ "n2^n" , then

"\pi_n(x)=\frac{i-1}{2^n}"

On the other hand,

"\pi_{n+1}(x)=\frac{2i-2}{2^n}" or "\frac{2i-1}{2^n}"


both of which are

"\ge \frac{i-1}{2^n}"


If x ∈ En then "\pi_n"(x) = n. But,

"f^{-1}([n,\infin])=f^{-1}(n,n+1)\bigcup f^{-1}([n+1,\infin])"

If "x\isin =f^{-1}(n,n+1)", then

"\pi_{n+1}(x)=\displaystyle{\sum_{i=n2^{n+1}+1}^{(n+1)2^{n+1}}}\frac{i-1}{2^{n+1}}\chi_{E_{n,i}}\ge \frac{n2^{n+1}}{2^{n+1}}=\pi_n(x)"


If x ∈ f−1([k + 1,∞]), then "\pi_{n+1}"(x) = k + 1 > k = "\pi_{n}" . thus, we have shown that for every x ∈ E,

"\pi_{n+1}(x)\ge \pi_{n}(x)"